Q7R.2 Suppose we have a box that emits electrons in a definite but unknown spin state | ψ). If we send electrons

media%2F9cb%2F9cbc19b0-d921-4134-a901-a5
Answer #1

Aluo 0-120 so we nud to find . a as b13 2-a dv3 204) СУ a = o-G061 母bz 0.7951 = 0.6064